How si the solution of
x-y=3
x+2y=6

Answer by mananth(16946)

1 x -1 y = 3 .............1
Total value
1 x + 2 y = 6 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 1
2 x -2 y = 6
1 x + 2 y = 6
Add the two equations
3 x = 12
/ 3
x = 4
plug value of x in (1)
1 x -1 y = 3
4 -1 y = 3
-1 y = 3 -4
-1 y = -1
y = 1